the length of a photograph is 8 cm more than its width. It has an area of 20cm2. Write an equation and solve to find the dimensions of the photograph
Answer by TimothyLamb(4379)

L = w + 8
Lw = 20
---
(w + 8)w = 20
ww + 8w - 20 = 0
---
the above quadratic equation is in standard form, with a=1, b=8 and c=-20
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 8 -20
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
w = 2
w = -10
---
the negative root doesn't fit the problem statement, so use the positive root:
---
answer:
w = 2 cm
L = 10 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php